OpenStudy (anonymous):

p(x)=x^4-6x^3+14x^2-24x+40 two roots are r_1= 2i and r_2= 3-i. Doing no polynomial or synthetic division, do the following: Find two more roots of p(x), and Write a complete factorization of p(x)

OpenStudy (blockcolder):

Complex roots and irrational roots always come in pairs. If a+bi is a factor of P(x), then a-bi is also a factor.

OpenStudy (anonymous):

the other two roots are complex conjugate of 2i and 3-i thus the other two roots are -2i and 3+i hence the complete factorisation can be written as (x-2i)(x+2i)(x-3+i)(x-3-i) =(x^2+4)((x-3)^2+1) =(x^2+4)(x^2 -6x +9+1) =(x^2+4)(x^2 -6x +10)

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

welcome dear